OpenStudy (anonymous):

what is the sin, cos, tan, csc, sec, and cot of 0 degrees?

jimthompson5910 (jim_thompson5910):

Do you have a unit circle with you?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

ok locate the angle 0 degrees what is the corresponding point on the unit circle?

OpenStudy (anonymous):

1,0 ?

jimthompson5910 (jim_thompson5910):

use parenthesis but yeah, (1,0)

jimthompson5910 (jim_thompson5910):

the x coordinate is equal to cos(theta) the y coordinate is equal to sin(theta)

jimthompson5910 (jim_thompson5910):

so this means `cos(0) = 1` and `sin(0) = 0`

OpenStudy (anonymous):

tan(0)= 0?

jimthompson5910 (jim_thompson5910):

to compute `tan(0)` you use this trig identity \[\Large \tan(\theta) = \frac{\sin(\theta)}{\cos(\theta)}\]

jimthompson5910 (jim_thompson5910):

yeah you'll find that tan(0) = sin(0)/cos(0) = 0/1 = 0

jimthompson5910 (jim_thompson5910):

for the other three, use the identities listed on page 2 under the "Reciprocal Identities" section http://tutorial.math.lamar.edu/pdf/Trig_Cheat_Sheet.pdf

jimthompson5910 (jim_thompson5910):

eg: csc(x) = 1/(sin(x))

OpenStudy (anonymous):

csc(0)= 0?

jimthompson5910 (jim_thompson5910):

what was sin(0) equal to again?

OpenStudy (anonymous):

0

jimthompson5910 (jim_thompson5910):

so csc(x) = 1/(sin(x)) csc(0) = 1/(sin(0)) csc(0) = 1/0 csc(0) = UND or NaN UND = undefined NaN = Not a Number

jimthompson5910 (jim_thompson5910):

1/0 is NOT the same as 0/1 0/1 = 0 1/0 is undefined because you can't have 0 in the denominator

OpenStudy (anonymous):

so i put undefined?

jimthompson5910 (jim_thompson5910):

yeah

jimthompson5910 (jim_thompson5910):

for csc(0)

OpenStudy (anonymous):

and for sec i also put undefined

jimthompson5910 (jim_thompson5910):

why? what was cos(0) ?

OpenStudy (anonymous):

1

jimthompson5910 (jim_thompson5910):

take the reciprocal of that to find sec(0)

jimthompson5910 (jim_thompson5910):

another way to view why csc(0) was undefined is that 0 doesn't have a reciprocal however, sec(0) is defined

OpenStudy (anonymous):

1?

jimthompson5910 (jim_thompson5910):

yeah sec(0) = 1

OpenStudy (anonymous):

cot(0)= undefined?

jimthompson5910 (jim_thompson5910):

why?

OpenStudy (anonymous):

tan(0)=0

jimthompson5910 (jim_thompson5910):

good, and taking the reciprocal of 0 leads to undefined

jimthompson5910 (jim_thompson5910):

so to summarize sin(0) = 0 cos(0) = 1 tan(0) = 0 csc(0) = und sec(0) = 1 cot(0) = und

OpenStudy (anonymous):

thanks.

jimthompson5910 (jim_thompson5910):

np